OpenStudy (anonymous):

1. Is the equation true, false, or open? 9p + 8 = 10p + 7 (1 point) Open; there is a variable. True; the expressions are the same for all values of the variables. False; the expressions are never the same. 2. Which value is a solution of the equation 5 – 4x = –3? (1 point) 0 2 –3 3. Which ordered pair is a solution of the equation y = x – 2? (1 point) (3, 1) (1, 3) (–1, 3) (3, –1) 4. Which ordered pair is a solution of the equation y = 5x? (1 point) (–2, 10) (–5, 25) (–3, 15) (–2, –10) 5. Which ordered pair is a solution of the equation y = –7x + 2? (1 point) (1, 2) (8, –

OpenStudy (anonymous):

1. Is the equation true, false, or open? 9p + 8 = 10p + 7 (1 point) Open; there is a variable. True; the expressions are the same for all values of the variables. False; the expressions are never the same. 2. Which value is a solution of the equation 5 – 4x = –3? (1 point) 0 2 –3 3. Which ordered pair is a solution of the equation y = x – 2? (1 point) (3, 1) (1, 3) (–1, 3) (3, –1) 4. Which ordered pair is a solution of the equation y = 5x? (1 point) (–2, 10) (–5, 25) (–3, 15) (–2, –10) 5. Which ordered pair is a solution of the equation y = –7x + 2? (1 point) (1, 2) (8, –54) (5, –35) (1, –7) 6. Nick and his cousin Sara have the same birthday, but Nick is eight years older than Sara. Let the variable x represent Nick’s age and y represent Sara’s age. Which graph represents the relationship between Nick’s age and Sara’s age? (1 point) 7. Solve the following equation. –3 = z + 4 (1 point) z = 1 z = –1 z = –7 z = 7 8. Solve the following equation. z – 4 = –8 (1 point) z = –2 z = 32 z = –4 z = –12 9. Solve the following equation. 6.4 = 4x (1 point) x = 2.4 x = 1.6 x = 4 x = 6.4 10. Solve the following equation. –9 = (1 point) r = 36 r = 13 r = r = –36 11. Solve the following equation. –7 = (1 point) z = z = z = z = 12. Which equation is an identity? (1 point) 8 – (5x + 2) = –5x – 6 7z + 10 – z = 8z – 2(z – 5) 8m – 4 = 5m + 8 – m 6y + 5 = 6y – 5 13. Which equation has no solution? (1 point) 7v + 2 = 8v – 3 3x – 5 = 3x + 8 – x 4y + 5 = 4y – 6 7z + 6 = –7z – 5 14. Solve the following equation. 5(z – 4) – z = 4z – 20 (1 point) 4 –4 infinitely many solutions no solution 15. Solve the following equation. 4 + 2y = 9 + 2y (1 point) 4 –4 infinitely many solutions no solution 16. Eric and Mark went to a ball game and visited the concession stand. They each got a hot dog, and Mark got a large soda and a bag of peanuts. Eric had a coupon for $3 off the cost. Mark paid the rest, which came to $11.50. Peanuts cost $4, and a hot dog costs $3.50. What was the cost for a large soda? (1 point) $7.50 $1.50 $4.00 $3.50 17. Solve the following equation. 6(z – 2) + 4 = 10 (1 point) z = 2 z = 3 z = –3 z = –4 18. Solve the following equation. 96 = –6(–4 – 3y) (1 point) y = 2 y = 4 y = –2 y = –4 19. Solve the following equation. –6p + 7 = 3(2p – 3) – 4(–10 + 4p) (1 point) p = 6 p = 5 p = 7 p = 12 20. WorkPad Note: Remember to show all of the steps that you use to solve the problem. Be sure to use the text box where the question mark (?) first appears to show your mathematical work. You can use the comments field to explain your work. Your teacher will review each step of your response to ensure you receive proper credit for your answer. Solve the following equation. u + 2 = 4 (2 points) 21. Solve the following equation. –5 = s – 3 (2 points) 22. Solve the following equation. t = –9 (2 points) 23. Solve the following equation. 6 = (2 points) 24. Solve the following equation. 13 + = –18 (2 points) Short Answer Note: Your teacher will grade your response to ensure you receive proper credit for your answer. 25. What is an estimate of the solution of the equation 6n + 3 = 2? Use a table. (5 points) 26. Solve the equation. 17 = –13 – 8x (2 points)

OpenStudy (anonymous):

@Love_Ranna

OpenStudy (anonymous):

@Ranaa

OpenStudy (anonymous):

@Love_Ranaa

OpenStudy (anonymous):

Missy you need to get better at Math because a seventh grader can do this.

OpenStudy (anonymous):

Well I'm sorry my learning disability inconveniences you

OpenStudy (cwrw238):

try to solve the first one. If you get a value for x it is open.

OpenStudy (anonymous):

Ok... That still doesn't help me I don't know how to do this my old school didn't teach me how to do any of this

OpenStudy (anonymous):

www.mathpapa.com

OpenStudy (arabpride):

Please post just one question per post, but if it's a test, \(\color{lightblue}{Open}\)\(\color{lightgreen}{study}\) users cannot give you answers. @Angel_wings

OpenStudy (arabpride):

And @Jamie_elizabeh Please do not 'advertise' for other sites on this one. Thanks.